OpenStudy (anonymous):

Simplify- a) log 25/log 0.2

OpenStudy (anonymous):

as in compute the number, or as in rewrite in a different from?

OpenStudy (anonymous):

\[\frac{ \log25 }{ \log 0.2 } = \frac{ \log 25 }{ \log \frac{ 1 }{ 5 } }\]

OpenStudy (anonymous):

\[\frac{ \log5^2 }{ \log 5^\left( -1) \right) }\]

OpenStudy (anonymous):

\[\frac{ 2\log5 }{ -1\log5 }\]

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

THANK YOU SO MUCH!!!!!! :D